A tennis ball is thrown vertically upward with an initial velocity of +8.0 m/s.


(a) what will the ball\'s speed be when it returns to its starting point?


(b) how long will the ball take to reach its starting point?

Grade:12

1 Answers

vikas askiitian expert
13 years ago

ans a) speed of ball will be equal to the speed of projection....

ans b)  applying v = u +at

 at the topmost point v=0

        t=u/a =8/10 sec                   (  accleration a =-g)

t is the time taken by the ball to move to highest point

 so total time of flight is 2t

   T=16/10 =8/5 sec